# RCCDA: Adaptive Model Updates in the Presence of Concept Drift under a Constrained Resource Budget

### v2

@@ 1 Introduction
In recent years, deep learning (DL) models have become popular in resource-constrained environments [liu2024mobilellm, NEURIPS2024_adb77ecc, 8373692, mehta2022mobilevit], e.g., on mobile devices, where computational power and memory are limited. These settings typically assume static data distributions during inference, in part due to resource limitations [tan2019efficientnet, pmlr-v202-castiglia23a, NEURIPS2023_dfee0949]. In practice, this poses a major challenge in the form of concept drift, which is a phenomenon where the underlying relationships in data shift over time, causing model performance to degrade [NEURIPS2022_46a12649, yao2022wildtime]. The computational demands of frequent updates that may be necessary to overcome rapid drifts present a significant challenge in resource-constrained settings. As a result, effectively managing model updates to adapt to such drift while adhering to resource constraints becomes a critical problem.
To address concept drift and mitigate model performance deterioration, researchers have proposed a variety of adaptive learning methods [bifet2007learning, 5719616, 5975223]. These strategies are often tailored to different operational contexts and environment characteristics [helli2024drift, chen2024classifier, zhang2023onenet, panchal2023flash]. While most are heuristic, validating their efficacy empirically [9427288, 8246541], certain techniques have offered convergence guarantees to give principled insights into algorithm design, e.g., [pmlr-v139-tahmasbi21a, NEURIPS2023_1fe6f635]. Even so, these theoretical analyses often rely on specific assumptions about the drift, or require prior knowledge about the environment that may not be readily available for decision-making.
@@
  • Optimal Policy Design: We formulate and solve an optimization problem to determine the optimal times for updating a model, minimizing the loss induced by concept drift while satisfying resource constraints. Leveraging the Lyapunov drift-plus-penalty framework [neely2010stochastic], we develop RCCDA: a threshold-based policy that triggers updates based on inference loss data and a virtual queue tracking resource expenditure. The resulting lightweight algorithm relies solely on historical loss data and a tunable drift threshold, making it efficient for real-time applications, while ensuring a greedy-optimal balance between model performance and computational overhead.
@@ 2 Related Work
Concept Drift. Many of the existing works on studying concept drift have placed an emphasis on characterizing or detecting drift [webb2016characterizing, agrahari2022concept, gama2004learning, dos2016fast, panda2023featureinterpretable]. For instance, windowing techniques, such as the Adaptive Windowing (ADWIN) algorithm [pmlr-v119-hinder20a], dynamically adjust the size of a data window to monitor statistical changes, flagging drift when significant deviations occur. Alternatively, observing performance indicators [harel2014concept, 8489364, baena2006early, cobb2022context, li2021detecting] involves tracking metrics like accuracy or error rates, where a sudden decline may signal that the model no longer aligns with the current data distribution. Additionally, machine learning models can be trained to detect drift by identifying shifts in feature distributions or relationships among variables [7372248, 9417164]. Detecting drift is a crucial step in drift adaptation, as it provides actionable insights into the optimal times to update or retrain the model with fresh data [pmlr-v139-tahmasbi21a, brzezinski2013reacting]. However, these methods do not consider operating under limited resource budget, a problem that is addressed by the policy introduced in our paper.
Domain Generalization. Domain generalization (DG) techniques address the related problem of training models to generalize well across multiple domains, aiming to improve robustness to unseen target distributions during inference [9782500, 9847099, wang2022generalizing, xie2023evolving]. Classical strategies include data/feature augmentation [zhoudomain, liuncertainty, li2021simple] or leveraging domain-invariant features [muandet2013domain, lu2022domaininvariant, NEURIPS2021_b0f2ad44]. More recent works have investigated temporal domain generalization [bai2023temporal, cai2024continuous, NEURIPS2023_459a911e] to adapt to time-varying distributions, as encountered during concept drift. In practice, however, these approaches demand significant resources, such as continuous data monitoring and frequent model updates, making them impractical for resource-constrained settings. Furthermore, these works have not explicitly considered adherence to resource constraints, which is one of the key motivations of our paper.
Continual Learning. As mitigating concept drift involves retraining the model at strategic intervals guided by a policy, the drift-adaptation problem shares similarities with continual learning (CL), where the goal is to enable models to learn from a sequence of tasks over time [10444954, NEURIPS2018_cee63112]. A major hurdle in CL is the problem of catastrophic forgetting, where adapting to new tasks causes the model to lose performance on previously learned ones [french1999catastrophic, kirkpatrick2017overcoming]. The problem studied in this paper prioritizes a different objective than CL: optimizing model performance on the current data distribution at each time step under resource constraints, i.e., immediate adaptability versus long-term memory retention. As such, our work is applicable to settings where the concept drift does not exhibit cyclic patterns.
@@ 3.2 Problem Formulation
For the sake of practicality and computational efficiency of our solution, we consider policies that are causal and distribution ignorant, meaning (i) the decision maker neither evaluates nor knows the data distributions $p_{0},p_{1},\dots p_{T-1}$ and (ii) the decision maker decides on the action at time $t$ based on the history of expected losses of the past inference batches of data $\mathcal{\xi}_{0},\mathcal{\xi}_{1},\ldots,\mathcal{\xi}_{t}$, which are available through running the inference task. This is an important distinction of our setting, as any further modifications to the action space would result in an increased task overhead. For example, trying to detect the concept drift requires performing operations on the dataset, which is highly resource consuming (see Appendix D for more details). Using the inference loss to evaluate the model performance is a common established strategy [he2014practical].
@@ 4 Policy Design
Idea and Overview. The central challenge of this work is adapting to changing conditions while respecting resource constraints. This issue spans beyond machine learning, appearing in areas where responsiveness and resource management must align, and requires strategies that balance performance and efficiency across dynamic systems. Existing solutions often rely on computationally expensive methods or fail to provide strict guarantees on resource usage and performance. To overcome these limitations, we leverage the Lyapunov drift-plus-penalty framework [neely2010stochastic], as it allows us to transform an intractable long-term optimization problem into a series of greedy, per-time-step decisions. The core intuition is to dynamically incorporate the long-term constraint into each per-time-step objective. This is achieved by defining a “penalty” term, which represents our immediate objective (based on minimizing the convergence bound), and a “drift” term, which measures how far the system deviates from its long-term resource budget. By optimizing a bound on this combined drift-plus-penalty expression at each time $t$, the agent is compelled to make greedy, online decisions that provably balance these two competing factors. This process creates an explicit, online trade-off between short-term performance and long-term budget adherence, allowing us to derive a principled, lightweight policy with provable guarantees.
@@ 5.2 Stability Analysis
Discussion. Theorem 5.3 guarantees that RCCDA satisfies the resource constraint in (5) as $T\to\infty$, provided the cumulative drift $\sum_{t=0}^{T-1}\sqrt{\delta(t)}$ grows sublinearly in $T$ and the update costs $\lambda(t)$ remain bounded, i.e., $\lambda(t)\sim\mathcal{O}(1)$. This reflects a key strength of the Lyapunov drift-plus-penalty framework [neely2010stochastic]. For finite $T$, the violation bound can be tightened toward zero by tuning V optimally. Additionally, under high but bounded drift rates ($\delta(t)\sim\mathcal{O}(1)$), the constraint violation decays as $\mathcal{O}(1/\sqrt{T})$, ensuring the policy adheres to the budget over long horizons without requiring drift to vanish.
@@ 6 Experiments
Datasets. To represent concept drift, we utilize four widely used domain generalization datasets: PACS [DBLP:journals/corr/abs-2007-14390], DigitsDG [10.1007/978-3-030-58517-4_33], and in the appendix OfficeHome [venkateswara2017deep], and MEMD-ABSA [cai2025memd]. The drift is simulated by introducing samples from different domains into the dataset. The model architectures utilized are detailed in Appendix E.2. Initially, we pretrain the models on data from a single source domain.
@@
  • Periodic: This deterministic policy updates the model after every time interval $\lambda/\bar{\lambda}$, i.e. $\pi_{\text{Periodic}}=1$ if $t=k\lceil\lambda/\bar{\lambda}\rceil,k\in\mathbb{N}\cup\{0\}$, and 0 otherwise. The policy is modified from [garcia2025concept].
  • Budget-Increase: The policy keeps track of an update budget available over time, then if it detects $n\in\mathbb{N}$ consecutive loss increases and if there is budget available, the model is updated. Based on modified trend-detection methods in [gama2014survey].
  • Budget-Threshold: The policy keeps track of an update budget and a window of losses $\mathcal{W}_{f}$. Then if $f_{t}\geq(1+\epsilon)\max_{f_{i}\in\mathcal{W}_{f}}f_{i}$ and if there is budget available, the model is updated. Based on a modified version of [bifet2007learning].
@@ Appendix B Convergence Proofs
    is valid because the policy’s action $\pi(t+1)$ is independent of the training data batch $\zeta_{t+1}$ conditioned on the past history, as $\pi(t+1)$ depends on inference batches $\xi_{t}$, independent of $\zeta_{t}$.
@@
The drift-induced loss term can related to the dynamic bound on the concept drift as follows:
@@ D.2 Overhead Analysis of the RCCDA Policy
To demonstrate the necessity for a lightweight policy that does not rely on explicit drift detection, we conducted an empirical analysis comparing RCCDA against two established drift detection methods: the Adaptive Windowing (ADWIN) [pmlr-v119-hinder20a] algorithm and the Kolmogorov-Smirnov (KS) [massey1951kolmogorov] test. Traditional drift detection methods like these typically require constant monitoring of the statistical properties of the entire incoming data stream (KS-Test) or sequentially processing error rates (ADWIN). In contrast, RCCDA relies solely on the model’s inference loss, which is a lightweight scalar value, often computed or estimated as part of the model’s primary task evaluation. The reliance on a pre-existing simple signal makes our approach inherently efficient.
@@ E.1 Datasets
  • PACS [li2017deeperbroaderartierdomain]: PACS consists of images from 4 domains: Photo, Art Painting, Cartoon, and Sketch, each with the 7 common categories: dog, elephant, giraffe, guitar, horse, house, and person. The default size of the images is $224\times 224$. We transform all images into $128\times 128$.
  • DigitsDG [10.1007/978-3-030-58517-4_33]: DigitsDG encompasses images of handwritten digits 0 to 9, sourced from 4 different domains: SVHN, SYN, MNIST, and MNIST-M. The image sizes are $32\times 32$.
  • OfficeHome [venkateswara2017deep]: OfficeHome consists of images from four domains: Art, Clipart, Product, and Real-World, each featuring 65 object categories commonly encountered in office and home environments, including chairs, desks, and household items. The image sizes are $224\times 224$.
  • MEMD-ABSA [cai2025memd]: MEMD-ABSA is a multi-domain textual dataset for Aspect-Based Sentiment Analysis (ABSA). It consists of approximately 20,000 sentences from 5 distinct domains: Books, Clothing, Hotel, Laptop and Restaurant. Each sentence annotated with one or more sentiment quadruples: aspect, category, opinion, and sentiment. The sentiment is then classified as positive, negative, or neutral.
@@ 20 Linear layer: 512 input features to 10 output classes
OfficeHome. For the OfficeHome dataset, we use OfficeHomeNet, a model based on a pretrained ResNet-18 architecture [he2015deepresiduallearningimage]. To adapt the model for the dataset’s complexity, we replace the original fully connected layer with a custom classification head. We fine-tune the model by training only the final convolutional block (layer4) and our custom head, keeping all earlier layers frozen. This results in a model with 8,541,761 trainable parameters out of 11,324,545 total. The custom head consists of:
@@ 4 Linear layer: 256 input features to 65 output classes
MEMD-ABSA. For the MEMD-ABSA dataset, we fine-tune a model based on the TinyBert architecture[jiao2019tinybert], called TinyBertForSentiment. The model comprises 4,386,307 parameters, and has the following structure:
@@ 6 Linear layer: 128 input features to 3 output classes
Each transformer encoder block is comprised of a 2-head self-attention mechanism with a hidden size of 128 and GELU activation functions. We tokenize input text using a pretrained BERT tokenizer that employs the WordPiece tokenization[schuster2012japanese]. All sequences are subsequently padded or truncated to a fixed length of 128 tokens.
@@ F.1 OfficeHome Dataset
In this section, we provide additional results on the OfficeHome [venkateswara2017deep] dataset, a domain generalization benchmark consisting of four domains: Art, Clipart, Product, and Real-World.
@@ F.2 MEMD-ABSA Dataset - Textual Modality
In this section, we provide additional results on the MEMD-ABSA dataset [cai2025memd], a multi-domain benchmark for textual Aspect-Based Sentiment Analysis (ABSA) with reviews from five domains: Books, Clothing, Hotel, Laptop and Restaurant. This experiment evaluates our method on a new, textual, modality, and explores its performance in a setting with less severe concept drift.
@@ Appendix G Code Documentation
  • flwr_datasets [DBLP:journals/corr/abs-2007-14390]

### v3

@@ 1 Introduction
In recent years, deep learning (DL) models have become popular in resource-constrained environments [33, 48, 41, 39], e.g., on mobile devices, where computational power and memory are limited. These settings typically assume static data distributions during inference, in part due to resource limitations [51, 11, 15]. In practice, this poses a major challenge in the form of concept drift, which is a phenomenon where the underlying relationships in data shift over time, causing model performance to degrade [46, 61]. The computational demands of frequent updates that may be necessary to overcome rapid drifts present a significant challenge in resource-constrained settings. As a result, effectively managing model updates to adapt to such drift while adhering to resource constraints becomes a critical problem.
To address concept drift and mitigate model performance deterioration, researchers have proposed a variety of adaptive learning methods [6, 40, 17]. These strategies are often tailored to different operational contexts and environment characteristics [25, 12, 62, 44]. While most are heuristic, validating their efficacy empirically [60, 49], certain techniques have offered convergence guarantees to give principled insights into algorithm design, e.g., [50, 38]. Even so, these theoretical analyses often rely on specific assumptions about the drift, or require prior knowledge about the environment that may not be readily available for decision-making.
@@
  • Optimal Policy Design: We formulate and solve an optimization problem to determine the optimal times for updating a model, minimizing the loss induced by concept drift while satisfying resource constraints. Leveraging the Lyapunov drift-plus-penalty framework [43], we develop RCCDA: a threshold-based policy that triggers updates based on inference loss data and a virtual queue tracking resource expenditure. The resulting lightweight algorithm relies solely on historical loss data and a tunable drift threshold, making it efficient for real-time applications, while ensuring a greedy-optimal balance between model performance and computational overhead.
@@ 2 Related Work
Concept Drift. Many of the existing works on studying concept drift have placed an emphasis on characterizing or detecting drift [56, 2, 19, 16, 45]. For instance, windowing techniques, such as the Adaptive Windowing (ADWIN) algorithm [26], dynamically adjust the size of a data window to monitor statistical changes, flagging drift when significant deviations occur. Alternatively, observing performance indicators [22, 14, 3, 13, 29] involves tracking metrics like accuracy or error rates, where a sudden decline may signal that the model no longer aligns with the current data distribution. Additionally, machine learning models can be trained to detect drift by identifying shifts in feature distributions or relationships among variables [36, 1]. Detecting drift is a crucial step in drift adaptation, as it provides actionable insights into the optimal times to update or retrain the model with fresh data [50, 7]. However, these methods do not consider operating under limited resource budget, a problem that is addressed by the policy introduced in our paper.
Domain Generalization. Domain generalization (DG) techniques address the related problem of training models to generalize well across multiple domains, aiming to improve robustness to unseen target distributions during inference [54, 63, 53, 58]. Classical strategies include data/feature augmentation [65, 32, 31] or leveraging domain-invariant features [42, 35, 8]. More recent works have investigated temporal domain generalization [4, 10, 57] to adapt to time-varying distributions, as encountered during concept drift. In practice, however, these approaches demand significant resources, such as continuous data monitoring and frequent model updates, making them impractical for resource-constrained settings. Furthermore, these works have not explicitly considered adherence to resource constraints, which is one of the key motivations of our paper.
Continual Learning. As mitigating concept drift involves retraining the model at strategic intervals guided by a policy, the drift-adaptation problem shares similarities with continual learning (CL), where the goal is to enable models to learn from a sequence of tasks over time [55, 59]. A major hurdle in CL is the problem of catastrophic forgetting, where adapting to new tasks causes the model to lose performance on previously learned ones [18, 28]. The problem studied in this paper prioritizes a different objective than CL: optimizing model performance on the current data distribution at each time step under resource constraints, i.e., immediate adaptability versus long-term memory retention. As such, our work is applicable to settings where the concept drift does not exhibit cyclic patterns.
@@ 3.2 Problem Formulation
For the sake of practicality and computational efficiency of our solution, we consider policies that are causal and distribution ignorant, meaning (i) the decision maker neither evaluates nor knows the data distributions $p_{0},p_{1},\dots p_{T-1}$ and (ii) the decision maker decides on the action at time $t$ based on the history of expected losses of the past inference batches of data $\mathcal{\xi}_{0},\mathcal{\xi}_{1},\ldots,\mathcal{\xi}_{t}$, which are available through running the inference task. This is an important distinction of our setting, as any further modifications to the action space would result in an increased task overhead. For example, trying to detect the concept drift requires performing operations on the dataset, which is highly resource consuming (see Appendix D for more details). Using the inference loss to evaluate the model performance is a common established strategy [24].
@@ 4 Policy Design
Idea and Overview. The central challenge of this work is adapting to changing conditions while respecting resource constraints. This issue spans beyond machine learning, appearing in areas where responsiveness and resource management must align, and requires strategies that balance performance and efficiency across dynamic systems. Existing solutions often rely on computationally expensive methods or fail to provide strict guarantees on resource usage and performance. To overcome these limitations, we leverage the Lyapunov drift-plus-penalty framework [43], as it allows us to transform an intractable long-term optimization problem into a series of greedy, per-time-step decisions. The core intuition is to dynamically incorporate the long-term constraint into each per-time-step objective. This is achieved by defining a “penalty” term, which represents our immediate objective (based on minimizing the convergence bound), and a “drift” term, which measures how far the system deviates from its long-term resource budget. By optimizing a bound on this combined drift-plus-penalty expression at each time $t$, the agent is compelled to make greedy, online decisions that provably balance these two competing factors. This process creates an explicit, online trade-off between short-term performance and long-term budget adherence, allowing us to derive a principled, lightweight policy with provable guarantees.
@@ 5.2 Stability Analysis
Discussion. Theorem 5.3 guarantees that RCCDA satisfies the resource constraint in (5) as $T\to\infty$, provided the cumulative drift $\sum_{t=0}^{T-1}\sqrt{\delta(t)}$ grows sublinearly in $T$ and the update costs $\lambda(t)$ remain bounded, i.e., $\lambda(t)\sim\mathcal{O}(1)$. This reflects a key strength of the Lyapunov drift-plus-penalty framework [43]. For finite $T$, the violation bound can be tightened toward zero by tuning V optimally. Additionally, under high but bounded drift rates ($\delta(t)\sim\mathcal{O}(1)$), the constraint violation decays as $\mathcal{O}(1/\sqrt{T})$, ensuring the policy adheres to the budget over long horizons without requiring drift to vanish.
@@ 6 Experiments
Datasets. To represent concept drift, we utilize four widely used domain generalization datasets: PACS [5], DigitsDG [64], and in the appendix OfficeHome [52], and MEMD-ABSA [9]. The drift is simulated by introducing samples from different domains into the dataset. The model architectures utilized are detailed in Appendix E.2. Initially, we pretrain the models on data from a single source domain.
@@
  • Periodic: This deterministic policy updates the model after every time interval $\lambda/\bar{\lambda}$, i.e. $\pi_{\text{Periodic}}=1$ if $t=k\lceil\lambda/\bar{\lambda}\rceil,k\in\mathbb{N}\cup\{0\}$, and 0 otherwise. The policy is modified from [21].
  • Budget-Increase: The policy keeps track of an update budget available over time, then if it detects $n\in\mathbb{N}$ consecutive loss increases and if there is budget available, the model is updated. Based on modified trend-detection methods in [20].
  • Budget-Threshold: The policy keeps track of an update budget and a window of losses $\mathcal{W}_{f}$. Then if $f_{t}\geq(1+\epsilon)\max_{f_{i}\in\mathcal{W}_{f}}f_{i}$ and if there is budget available, the model is updated. Based on a modified version of [6].
@@ Appendix B Convergence Proofs
    is valid because the policy’s action $\pi(t+1)$ is independent of the training data batch $\zeta_{t+1}$ conditioned on the past history.
@@
We can relate the drift-induced loss term with the dynamic bound on the concept drift as follows:
@@ D.2 Overhead Analysis of the RCCDA Policy
To demonstrate the necessity for a lightweight policy that does not rely on explicit drift detection, we conducted an empirical analysis comparing RCCDA against two established drift detection methods: the Adaptive Windowing (ADWIN) [26] algorithm and the Kolmogorov-Smirnov (KS) [37] test. Traditional drift detection methods like these typically require constant monitoring of the statistical properties of the entire incoming data stream (KS-Test) or sequentially processing error rates (ADWIN). In contrast, RCCDA relies solely on the model’s inference loss, which is a lightweight scalar value, often computed or estimated as part of the model’s primary task evaluation. The reliance on a pre-existing simple signal makes our approach inherently efficient.
@@ E.1 Datasets
  • PACS [30]: PACS consists of images from 4 domains: Photo, Art Painting, Cartoon, and Sketch, each with the 7 common categories: dog, elephant, giraffe, guitar, horse, house, and person. The default size of the images is $224\times 224$. We transform all images into $128\times 128$.
  • DigitsDG [64]: DigitsDG encompasses images of handwritten digits 0 to 9, sourced from 4 different domains: SVHN, SYN, MNIST, and MNIST-M. The image sizes are $32\times 32$.
  • OfficeHome [52]: OfficeHome consists of images from four domains: Art, Clipart, Product, and Real-World, each featuring 65 object categories commonly encountered in office and home environments, including chairs, desks, and household items. The image sizes are $224\times 224$.
  • MEMD-ABSA [9]: MEMD-ABSA is a multi-domain textual dataset for Aspect-Based Sentiment Analysis (ABSA). It consists of approximately 20,000 sentences from 5 distinct domains: Books, Clothing, Hotel, Laptop and Restaurant. Each sentence annotated with one or more sentiment quadruples: aspect, category, opinion, and sentiment. The sentiment is then classified as positive, negative, or neutral.
@@ 20 Linear layer: 512 input features to 10 output classes
OfficeHome. For the OfficeHome dataset, we use OfficeHomeNet, a model based on a pretrained ResNet-18 architecture [23]. To adapt the model for the dataset’s complexity, we replace the original fully connected layer with a custom classification head. We fine-tune the model by training only the final convolutional block (layer4) and our custom head, keeping all earlier layers frozen. This results in a model with 8,541,761 trainable parameters out of 11,324,545 total. The custom head consists of:
@@ 4 Linear layer: 256 input features to 65 output classes
MEMD-ABSA. For the MEMD-ABSA dataset, we fine-tune a model based on the TinyBert architecture[27], called TinyBertForSentiment. The model comprises 4,386,307 parameters, and has the following structure:
@@ 6 Linear layer: 128 input features to 3 output classes
Each transformer encoder block is comprised of a 2-head self-attention mechanism with a hidden size of 128 and GELU activation functions. We tokenize input text using a pretrained BERT tokenizer that employs the WordPiece tokenization[47]. All sequences are subsequently padded or truncated to a fixed length of 128 tokens.
@@ F.1 OfficeHome Dataset
In this section, we provide additional results on the OfficeHome [52] dataset, a domain generalization benchmark consisting of four domains: Art, Clipart, Product, and Real-World.
@@ F.2 MEMD-ABSA Dataset - Textual Modality
In this section, we provide additional results on the MEMD-ABSA dataset [9], a multi-domain benchmark for textual Aspect-Based Sentiment Analysis (ABSA) with reviews from five domains: Books, Clothing, Hotel, Laptop and Restaurant. This experiment evaluates our method on a new, textual, modality, and explores its performance in a setting with less severe concept drift.
@@ Appendix G Code Documentation
  • flwr_datasets [5]